OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f of x equals two divided by x and g of x equals two divided by x.

OpenStudy (anonymous):

f(g(x)) means that every x value you see in f(x), you would replace it with what g(x) equals.

OpenStudy (aum):

\[ f(x) = \frac 2x \\ g(x) = \frac 2x \\ f(g(x)) = f\left(\frac 2x\right) = ? \]